OpenStudy (anonymous):

MEDAL FOR HELP _Medal for help :) A student scored 83 and 91 on her first two quizzes. Write and solve a compound inequality to find the possible values for a third quiz score that would give her an average between 85 and 90 inclusive Look below for choices ty :)

OpenStudy (dbzfan836):

choices?

OpenStudy (dbzfan836):

what are they???

OpenStudy (anonymous):

@Dbzfan836 Sorry it wouldn't upload

OpenStudy (anonymous):

@Cosmichaotic

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

The average of the three quizzes would be \(\dfrac{83+91+n}{3}\), so the inequality for the average (in order for the final average to fall in the desired range) would be \[85\le\dfrac{83+91+n}{3}\le90\]

OpenStudy (anonymous):

So it's A? @SithsAndGiggles

OpenStudy (anonymous):

Yes, that's the only option that agrees.

OpenStudy (anonymous):

TY For breaking it down :)

OpenStudy (anonymous):

yw